OpenStudy (anonymous):

please help me! roland had already covered 14km when charlemagne headed out to catch him. Charlemagne's rate was 11kph, whereas Roland's rate was 7 1/2kph. How long did it take Charlemagne to catch Roland?

OpenStudy (whpalmer4):

Charlemagne's speed is 11 km/h, Roland's is 7.5 km/h. That means that Charlemagne gains (11 - 7.5) km/h = 3.5 km/h on Roland. If Roland has a head start of 14 km, we just need to divide 14 km by 3.5 km/h to find out how many hours it takes to catch up. Another way of looking at this is that Charlemagne's position at time t is 11t, and Roland's position is 14 + 7.5t. We set the two of them equal to each other and solve for t. 11t = 14 + 7.5t 3.5t = 14 t = 14/3.5 at the time they have both traveled the same distance.

OpenStudy (anonymous):

so the answer is 4h?

OpenStudy (whpalmer4):

Yep, 14/3.5 = 4 in my experience, your mileage may vary :-)